explain how to figure out how many chances for president, vice president, and secretary if there are 20 students
Answer by jim_thompson5910(35256)

There are 20 choices for the first slot

20-1 = 19 choices for the second slot

and 20-2 = 18 choices for the third


So there are 20*19*18 = 6840 possible permutations.